A buffer that contains 0.27 M of a base, B and 0.44 M of its conjugate acid BH+, has a pH of 9.05. What is the pH after 0.03 mol of NaOH are added to 0.79 L of the solution?
Answer #1

Anstoeri- Given, Buffeo Solution contains. Base - 0 WFM. conjugate acid = 0.44M :PH: 9.05 PH aftem:0.03 mol Of NaOH. Solutionpot = pkb + log [conjugare acid)/(base] Wiikit pOH = 14-PH = 149.05 - 4.95 Substitute the values in eo 4.958 PKB4109 [0.4410:hlikit PH = 14-POH =14-4.855 29.14 n. The pH after 0.03 mol of Naoh arel I added to 0.79L of solution is 9.14.